OpenStudy (anonymous):

MEDALS PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!!!!! Casey is completing an experiment that requires 300 ml of potassium chloride (PCl) at a concentration of 0.2 M. But Casey's lab only stores potassium chloride at concentrations of 0.1 M and 0.5 M. How much 0.1 M PCl will Casey have to mix with 0.5 M PCl to make 300 ml of 0.2 M PCl?

OpenStudy (anonymous):

can u help me plz

OpenStudy (amoodarya):

x from .1 m y from .5 m x(.1)+y(.5)=(x+y)(.2) x+y=300 ml